# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### 18 Dec 2023 v2

@@ A.1.1 Add Gaussian Noise to Original Images
where $\boldsymbol{I}_{k}$ is a $k\times k$ identity matrix and $\boldsymbol{1}_{k}$ is a all ones $k\times k$ matrix. We use $\boldsymbol{U}$ to represent $\mathrm{diag}(\sigma^{2}_{X_{1}},...,\sigma^{2}_{X_{k}})\boldsymbol{I}_{k}$, and $\boldsymbol{u}$ to represent a all ones vector $[1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula [50] and Woodbury Formula [63], we can obtain the inverse of $\boldsymbol{{\Sigma}_{X+\epsilon}}$ as:
@@
$\begin{split}\boldsymbol{{\Sigma}_{X+\epsilon}^{-1}}=&(\boldsymbol{U}+\sigma^{2}_{\epsilon}\boldsymbol{uu}^{T})^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\boldsymbol{u}}\boldsymbol{U}^{-1}\boldsymbol{u}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}}\boldsymbol{U}^{-1}\boldsymbol{1}_{k}\boldsymbol{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{X_{1}}}-\frac{1}{\sigma^{4}_{X_{1}}}&-\frac{1}{\sigma^{2}_{X_{1}}\sigma^{2}_{X_{2}}}&...&-\frac{1}{\sigma^{2}_{X_{1}}\sigma^{2}_{X_{k-1}}}&-\frac{1}{\sigma^{2}_{X_{1}}\sigma^{2}_{X_{k}}}\\ -\frac{1}{\sigma^{2}_{X_{2}}\sigma^{2}_{X_{1}}}&\frac{1}{\lambda\sigma^{2}_{X_{2}}}-\frac{1}{\sigma^{4}_{X_{2}}}&...&-\frac{1}{\sigma^{2}_{X_{2}}\sigma^{2}_{X_{k-1}}}&-\frac{1}{\sigma^{2}_{X_{2}}\sigma^{2}_{X_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{X_{k-1}}\sigma^{2}_{X_{1}}}&-\frac{1}{\sigma^{2}_{X_{k-1}}\sigma^{2}_{X_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{X_{k-1}}}-\frac{1}{\sigma^{4}_{X_{k-1}}}&-\frac{1}{\sigma^{2}_{X_{k-1}}\sigma^{2}_{X_{k}}}\\ -\frac{1}{\sigma^{2}_{X_{k}}\sigma^{2}_{X_{1}}}&-\frac{1}{\sigma^{2}_{X_{k}}\sigma^{2}_{X_{2}}}&...&-\frac{1}{\sigma^{2}_{X_{k}}\sigma^{2}_{X_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{X_{k}}}-\frac{1}{\sigma^{4}_{X_{k}}}\\ \end{bmatrix}\end{split}$ (47)
@@
where $\boldsymbol{U}^{-1}=\mathrm{diag}((\sigma^{2}_{X_{1}})^{-1},...,(\sigma^{2}_{X_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}}$.
@@
Therefore, substitute Equation 47 into ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}$, we can obtain:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\sigma^{2}_{Y_{k}}\end{bmatrix}-\begin{bmatrix}\mathrm{cov}(Y_{1},X_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},X_{k}+\epsilon)\end{bmatrix}\boldsymbol{{\Sigma}_{X+\epsilon}^{-1}}\begin{bmatrix}\mathrm{cov}(Y_{1},X_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},X_{k}+\epsilon)\end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\mathrm{cov}^{2}(Y_{1},X_{1}+\epsilon)(\frac{1}{\sigma^{2}_{X_{1}}}-\frac{\lambda}{\sigma^{4}_{X_{1}}})&...&\mathrm{cov}(Y_{1},X_{1}+\epsilon)\mathrm{cov}(Y_{k},X_{k}+\epsilon)\frac{\lambda}{\sigma^{2}_{X_{1}}\sigma^{2}_{X_{k}}}\\ \vdots&&\vdots\\ \mathrm{cov}(Y_{k},X_{k}+\epsilon)\mathrm{cov}(Y_{1},X_{1}+\epsilon)\frac{\lambda}{\sigma^{2}_{X_{k}}\sigma^{2}_{X_{1}}}&...&\sigma^{2}_{Y_{k}}-\mathrm{cov}^{2}(Y_{k},X_{k}+\epsilon)(\frac{1}{\sigma^{2}_{X_{k}}}-\frac{\lambda}{\sigma^{4}_{X_{k}}})\\ \end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\frac{1}{\sigma^{2}_{X_{1}}}\mathrm{cov}^{2}(Y_{1},X_{1})&&\\ &\ddots&\\ &&\sigma^{2}_{Y_{k}}-\frac{1}{\sigma^{2}_{X_{k}}}\mathrm{cov}^{2}(Y_{k},X_{k})\\ \end{bmatrix}+\lambda\begin{bmatrix}\frac{1}{\sigma^{4}_{X_{1}}}\mathrm{cov}^{2}(Y_{1},X_{1})&...&\frac{1}{\sigma^{2}_{X_{1}}\sigma^{2}_{X_{k}}}\mathrm{cov}(Y_{1},X_{1})\mathrm{cov}(Y_{k},X_{k})\\ \vdots&&\vdots\\ \frac{1}{\sigma^{2}_{X_{k}}\sigma^{2}_{X_{1}}}\mathrm{cov}(Y_{k},X_{k})\mathrm{cov}(Y_{1},X_{1})&...&\frac{1}{\sigma^{4}_{X_{k}}}\mathrm{cov}^{2}(Y_{k},X_{k})\\ \end{bmatrix}\right|\end{split}$
@@
We use the notation $\boldsymbol{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{X_{1}}}\mathrm{cov}(Y_{1},X_{1})&\cdots&\frac{1}{\sigma^{2}_{X_{k}}}\mathrm{cov}(Y_{k},X_{k})\end{bmatrix}^{T}$, and $\boldsymbol{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{X_{1}}}\mathrm{cov}^{2}(Y_{1},X_{1}),\cdots,\frac{1}{\sigma^{2}_{X_{k}}}\mathrm{cov}^{2}(Y_{k},X_{k}))$. And utilize the rule of determinants of sums [35], then we have:
$\begin{split}|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|=&|(\boldsymbol{\Sigma_{Y}-V})+\lambda\boldsymbol{v}\boldsymbol{v}^{T}|\\ =&|\boldsymbol{\Sigma_{Y}-V}|+\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})^{*}}\boldsymbol{v}\end{split}$ (49)
@@
where $(\boldsymbol{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\boldsymbol{\Sigma_{Y}-V})$. For simplicity, we can rewrite ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}$ as:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})+\Omega\end{split}$ (50)
@@
$\begin{split}\Omega=\lambda\begin{bmatrix}\frac{1}{\sigma^{2}_{X_{1}}}\mathrm{cov}(Y_{1},X_{1})&\cdots&\frac{1}{\sigma^{2}_{X_{k}}}\mathrm{cov}(Y_{k},X_{k})\end{bmatrix}\begin{bmatrix}V_{11}&&\\ &\ddots&\\ &&V_{kk}\end{bmatrix}\begin{bmatrix}\frac{1}{\sigma^{2}_{X_{1}}}\mathrm{cov}(Y_{1},X_{1})\\ \vdots\\ \frac{1}{\sigma^{2}_{X_{k}}}\mathrm{cov}(Y_{k},X_{k})\end{bmatrix}\end{split}$ (51)
@@
$V_{ii}=\prod_{j=1,j\neq i}^{k}\left[\sigma^{2}_{Y_{j}}-\frac{1}{\sigma^{2}_{X_{j}}}\mathrm{cov}^{2}(X_{j},Y_{j})\right]$ (52)
@@
$\begin{split}\Omega=&\lambda\sum_{i=1}^{k}\frac{\frac{1}{\sigma^{4}_{X_{i}}}\mathrm{cov}^{2}(Y_{i},X_{i})\prod_{j=1}^{k}(\sigma^{2}_{Y_{j}}-\mathrm{cov}^{2}(Y_{j},X_{j})\frac{1}{\sigma^{2}_{X_{j}}})}{(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})}\\ =&\lambda\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}\\ \end{split}$ (53)
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))})\end{split}$ (54)
Accordingly, $|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|$ is:
$|\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{X_{i}}}\mathrm{cov}^{2}(X_{i},Y_{i}))$ (55)
As a result, $\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}$ is expressed as:

### 02 Feb 2024 v3

@@ B.1 Inject Gaussian Noise in Latent Space
where $\boldsymbol{I}_{k}$ is a $k\times k$ identity matrix and $\boldsymbol{1}_{k}$ is a all ones $k\times k$ matrix. We use $\boldsymbol{U}$ to represent $\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}$, and $\boldsymbol{u}$ to represent a all ones vector $[1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula [51] and Woodbury Formula [64], we can obtain the inverse of $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ as:
@@
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}^{-1}}=&(\boldsymbol{U}+\sigma^{2}_{\epsilon}\boldsymbol{uu}^{T})^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\boldsymbol{u}}\boldsymbol{U}^{-1}\boldsymbol{u}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}\boldsymbol{U}^{-1}\boldsymbol{1}_{k}\boldsymbol{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{Z_{1}}}-\frac{1}{\sigma^{4}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{2}}}-\frac{1}{\sigma^{4}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{Z_{k-1}}}-\frac{1}{\sigma^{4}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{k}}}-\frac{1}{\sigma^{4}_{Z_{k}}}\\ \end{bmatrix}\end{split}$ (51)
@@
where $\boldsymbol{U}^{-1}=\mathrm{diag}((\sigma^{2}_{Z_{1}})^{-1},...,(\sigma^{2}_{Z_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$.
@@
Therefore, substitute Equation 51 into ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$, we can obtain:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\sigma^{2}_{Y_{k}}\end{bmatrix}-\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\boldsymbol{{\Sigma}_{Z+\epsilon}^{-1}}\begin{bmatrix}\mathrm{cov}(Y_{1},Z_{1}+\epsilon)&...&0\\ \vdots&\ddots&\vdots\\ 0&...&\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\mathrm{cov}^{2}(Y_{1},Z_{1}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{1}}}-\frac{\lambda}{\sigma^{4}_{Z_{1}}})&...&\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\mathrm{cov}(Y_{k},Z_{k}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ \vdots&&\vdots\\ \mathrm{cov}(Y_{k},Z_{k}+\epsilon)\mathrm{cov}(Y_{1},Z_{1}+\epsilon)\frac{\lambda}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&...&\sigma^{2}_{Y_{k}}-\mathrm{cov}^{2}(Y_{k},Z_{k}+\epsilon)(\frac{1}{\sigma^{2}_{Z_{k}}}-\frac{\lambda}{\sigma^{4}_{Z_{k}}})\\ \end{bmatrix}\right|\\ =&\left|\begin{bmatrix}\sigma^{2}_{Y_{1}}-\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&&\\ &\ddots&\\ &&\sigma^{2}_{Y_{k}}-\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}+\lambda\begin{bmatrix}\frac{1}{\sigma^{4}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{1},Z_{1})\mathrm{cov}(Y_{k},Z_{k})\\ \vdots&&\vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{k},Z_{k})\mathrm{cov}(Y_{1},Z_{1})&...&\frac{1}{\sigma^{4}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k})\\ \end{bmatrix}\right|\end{split}$
@@
We use the notation $\boldsymbol{v}=\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}^{T}$, and $\boldsymbol{V}=\mathrm{diag}(\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}^{2}(Y_{1},Z_{1}),\cdots,\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}^{2}(Y_{k},Z_{k}))$. And utilize the rule of determinants of sums [36], then we have:
$\begin{split}|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|=&|(\boldsymbol{\Sigma_{Y}-V})+\lambda\boldsymbol{v}\boldsymbol{v}^{T}|\\ =&|\boldsymbol{\Sigma_{Y}-V}|+\lambda\boldsymbol{v}^{T}\boldsymbol{(\boldsymbol{\Sigma_{Y}-V})^{*}}\boldsymbol{v}\end{split}$ (53)
@@
where $(\boldsymbol{\Sigma_{Y}-V})^{*}$ is the adjoint of the matrix $(\boldsymbol{\Sigma_{Y}-V})$. For simplicity, we can rewrite ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$ as:
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})+\Omega\end{split}$ (54)
@@
$\begin{split}\Omega=\lambda\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})&\cdots&\frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\begin{bmatrix}V_{11}&&\\ &\ddots&\\ &&V_{kk}\end{bmatrix}\begin{bmatrix}\frac{1}{\sigma^{2}_{Z_{1}}}\mathrm{cov}(Y_{1},Z_{1})\\ \vdots\\ \frac{1}{\sigma^{2}_{Z_{k}}}\mathrm{cov}(Y_{k},Z_{k})\end{bmatrix}\end{split}$ (55)
@@
$V_{ii}=\prod_{j=1,j\neq i}^{k}\left[\sigma^{2}_{Y_{j}}-\frac{1}{\sigma^{2}_{Z_{j}}}\mathrm{cov}^{2}(Z_{j},Y_{j})\right]$ (56)
@@
$\begin{split}\Omega=&\lambda\sum_{i=1}^{k}\frac{\frac{1}{\sigma^{4}_{Z_{i}}}\mathrm{cov}^{2}(Y_{i},Z_{i})\prod_{j=1}^{k}(\sigma^{2}_{Y_{j}}-\mathrm{cov}^{2}(Y_{j},Z_{j})\frac{1}{\sigma^{2}_{Z_{j}}})}{(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})}\\ =&\lambda\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}\\ \end{split}$ (57)
@@
$\begin{split}&{|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}\\ =&\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})\end{split}$ (58)
Accordingly, $|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|$ is:
$|\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|=\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))$ (59)
As a result, $\frac{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}{|\boldsymbol{\Sigma_{Y|Z}}|}$ is expressed as: